OpenStudy (loser66):

Teach me, please Find all solutions to the congruence 55x = 35 (mod 75)

OpenStudy (loser66):

@ganeshie8

ganeshie8 (ganeshie8):

gcd(75,55) = 5 divide 5 through out

ganeshie8 (ganeshie8):

11x = 7 (mod 15)

OpenStudy (loser66):

yes, my bad

ganeshie8 (ganeshie8):

and why are we allowed to divide like that ?

ganeshie8 (ganeshie8):

you might be knowing that already, so lets move on and find the x in mod 15

OpenStudy (loser66):

Because 75,35 and 55 all are multiple of 5, so reduce to the lowest modulo, right?

ganeshie8 (ganeshie8):

yes 11x = 7 (mod 15) Notice that x = 2 gives : 11*2 = 7 (mod 15) 22 = 7 (mod 15) 15 = 0 (mod 15) which is true, so `x = 2(mod 15)` is a solution

ganeshie8 (ganeshie8):

but you want to express the solution in mod 75 right ?

OpenStudy (loser66):

yes,

OpenStudy (loser66):

but how to get 2 there? just try all remainder in Z_15?

ganeshie8 (ganeshie8):

thats one way when the numbers are small

ganeshie8 (ganeshie8):

the general method is to use euclidean algorithm

ganeshie8 (ganeshie8):

heard of it before ?

OpenStudy (loser66):

2years ago, once I saw it. Not sure I can remember or not, remind me, please

ganeshie8 (ganeshie8):

11x = 7 (mod 15) is equivalent to 11x = 15y + 7 for some y, right ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

which is same as 11x - 15y = 7

ganeshie8 (ganeshie8):

Notice that our goal here is to find the solutions in `integer values of x and y`

ganeshie8 (ganeshie8):

so this is a linear diophantine equation

ganeshie8 (ganeshie8):

the method is as follows : express gcd(11, 15) =1 as linear combination of 11 and 15 and multiply both sides by 7

ganeshie8 (ganeshie8):

we can express gcd(a,b) as linear combination of a,b : ax + by = gcd(a,b) u knew this right ?

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

good, then its easy for you : we digress from our goal a bit and solve 11x+15y=1 first : 15 = 11*1 + 4 11 = 4*2 + 3 4 = 3*1 + 1 3 = 1*3 + 0 so the gcd(11, 15) is 1

OpenStudy (loser66):

3(15)-4(11)

ganeshie8 (ganeshie8):

Now go in reverse : 1 = 4 - 3(1) = 4 - (11 - 4*2) = 15-11*1 - (11 - (15-11*1)*2) = 15 - 11 - 11 + 2(15-11) = 15(3) + 11(-4)

ganeshie8 (ganeshie8):

1 = 15(3) + 11(-4) multiply both sides by 7, you get : 7 = 15(21) + 11(-28) y = 21, x = -28

ganeshie8 (ganeshie8):

11x = 7 (mod 15) solution is x = -28 (mod 15)

ganeshie8 (ganeshie8):

let me know if something is not clear

OpenStudy (loser66):

I got it so far.

ganeshie8 (ganeshie8):

we're done, the solution is x = -28(mod 15) which is same as x= 2(mod 15)

ganeshie8 (ganeshie8):

you can express it in mod 75 you would get 75/15 = 5 incongruent solutions in mod 75

ganeshie8 (ganeshie8):

x = 2 (mod 15) is same as x = 2 + 15t for some t, right ?

ganeshie8 (ganeshie8):

plugin t = 0,1,2,3,4

OpenStudy (loser66):

yes

ganeshie8 (ganeshie8):

bascially you want all the x values that are less than 75 when you're in mod75

OpenStudy (loser66):

and check which values give us 55x =35 (mod 75) right?

ganeshie8 (ganeshie8):

No, we're done. all of those 5 values satisfy above ^^

ganeshie8 (ganeshie8):

x = 2 + 15t t = 0 : x = 2 t = 1 : x = 17 t = 2 : x = 32 t = 3 : x = 47 t = 4 : x = 62 so the incongruent solutions to 55x =35 (mod 75) are : {2, 17, 32, 47, 62}

OpenStudy (loser66):

Ok, If I plug t =1, then x =17 so that 55(17) = 935 and 935 = 12*(75) +35 so that 935 = 35 (mod 75)

OpenStudy (loser66):

Got it:))))))))).Thanks a lot

ganeshie8 (ganeshie8):

np :)